Section 08  Analysis

Section 08: Analysis

Invited Lectures

Wednesday, 19 August  MA 001

16:15 – 17:00 Michael T. Lacey, Georgia Institute of Technology, Atlanta, USA On the Bilinear Hilbert Transform (I: p.99) Chair: Elias M. Stein

17:15 – 18:00 Kristian Seip, Norwegian University of Science and Technology, Trondheim, Norway Developments from Nonharmonic Fourier Series (I: p.105) Chair: Klaus D. Bierstedt

Thursday, 20 August  MA 001

14:00 – 14:45 Sergey Ivanovich Pinchuk, Indiana University, Bloomington, USA and Chelyabinsk State Univ. of Technology, Chelyabinsk, Russia Reflection Principle in Higher Dimensions (I: p.104) Chair: Jürgen Leiterer

15:00 – 15:45 Nigel David Higson, Pennsylvania State University, University Park, USA The Baum-Connes Conjecture (I: p.98) Chair: Joachim Cuntz

Friday, 21 August  MA 001

16:15 – 17:00 Michael Christ, University of California, Berkeley, USA Singularity and Regularity — Local and Global (I: p.97) Chair: Daryl Geller

17:15 – 18:00 Hart F. Smith, University of Washington, Seattle, USA Wave Equations with Low Regularity Coefficients (I: p.107) Chair: Gerald B. Folland

Saturday, 22 August  MA 001

14:00 – 14:45 Kari Astala, University of Jyväskylä, Jyväskylä, Finland Analytic aspects of quasiconformality (I: p.97) Chair: Frederick W. Gehring

15:00 – 15:45 Detlef H. Müller, Universität Kiel, Kiel, Germany L^{p}-functional calculus on Lie groups and estimates for wave equations (I: p.102) Chair: N.N.

Monday, 24 August  MA 001

16:15 – 17:00 Stephen Wainger, University of Wisconsin, Madison, USA Discrete Analogues of Singular and Maximal Radon Transforms (I: p.109) Chair: Detlef H. Müller

17:15 – 18:00 Vitali Milman, Tel Aviv University, Tel Aviv, Israel Randomness and Pattern in Convex Geometric Analysis (I: p.101) Chair: Hermann König

Analysis  Section 08

Tuesday, 25 August  MA 001

14:00 – 14:45 Nicole Tomczak-Jaegermann, University of Alberta, Edmonton, Canada From finite- to infinite-dimensional phenomena in Geometric Functional Analysis, on local and asymptotic levels (I: p.108) Chair: W. Timothy Gowers

Wednesday, 26 August  MA 001

16:15 – 17:00 Pertti Mattila, University of Jyväskylä, Jyväskylä, Finland Rectifiability, analytic capacity, and singular integrals (I: p.100) Chair: Srishti (S.D.) Chatterji

17:15 – 18:00 Stefan Müller, Max-Planck-Inst. für Math. in den Naturwissenschaften, Leipzig, Germany Unexpected solutions of first and second order partial differential equations (I: p.104) Chair: John M. Ball

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Function Spaces and Operators  MA 041

Chair: Jose Bonet

14:00 – 14:15 Silke Holtmanns, Universität-Gesamthochschule, Paderborn, Germany An operator representation for weighted spaces of vector valued holomorphic functions (p.130)

14:15 – 14:30 Klaus D. Bierstedt, Universität-Gesamthochschule, Paderborn, Germany An operator representation for weighted inductive limits of spaces of vector valued holomorphic functions (p.117)

14:30 – 14:45 Jose Bonet, Universidad Politécnica de Valencia, Valencia, Spain Hypercyclic operators on non-normable Fréchet spaces (p.118)

Iteration, Composition, Clifford Structure  MA 309

Chair: Mario Bonk

14:00 – 14:15 Chung-Cun Yang, The Hong Kong University of Sci. & Tech., Hong Kong, China, P.R. of Some recent results on factorization of meromorphic functions (p.165)

14:15 – 14:30 Eugen J. Ionascu, Dept. of Math., University of Georgia, Athens, GA, USA Wandering sets for a class of Borel isomorphisms of [0, 1) (joint work with Dr. Edward Azoff, UGA, Athens, USA) (p.131)

14:30 – 14:45 Enrique Ramirez-De-Arellano, Dept. of Mathematics — Cinvestav, Mexico City, Mexico Almost-Clifford structures on manifolds (p.150)

Fractional Differentiation  MA 043

Chair: A. Eduardo Gatto

14:00 – 14:15 Evelyn Buckwar, FB Mathematik und Informatik, FU, Berlin, Germany Application of the Lie group of scaling transformations to the fractional diffusion-wave equation

(p.119)

14:15 – 14:30 Yuri Luchko, Dept. of Mathematics and Computer Science, Freie Universität, Berlin, Germany Lie group analysis of a partial differential equation of fractional order (p.140)

14:30 – 14:45 A. Eduardo Gatto, Depaul University, Chicago, USA Fractional Calculus, Sobolev Spaces, and Related Rules on Spaces of Homogeneous Type (p.124)

C*-Algebras and Modules  MA 041

Chair: Alexander Ya. Helemskii

15:00 – 15:15 Michael J. Gruber, Humboldt–Universität, Berlin, Germany Non-commutative Bloch theory (p.128)

15:15 – 15:30 Evgenij V. Troitsky, Moscow State University, Moscow, Russia Kuiper and Dixmier-Douady theorems for C^{∗}-Hilbert modules and K-theory (p.160)

15:30 – 15:45 Alexander Ya. Helemskii, Moscow State University, Moscow, Russia Description of projective Hilbert modules over C^{∗}-algebras (p.129)

Spaces of Analytic Functions  MA 309

Chair: Peter Duren

15:00 – 15:15 Mitsuo Morimoto, International Christian University, Mitaka-City, Japan Reproducing kernels related to the complex sphere (p.142)

15:15 – 15:30 Romeo Meštrović, Univerzitet Crne Gore, Kotor, Yugoslavia Topologies on Some F -algebras of Holomorphic Functions (p.141)

15:30 – 15:45 William E. Hornor, University of Southern Mississippi, Hattiesburg, Mississippi, USA Properties of isometry-inducing maps of the unit disc (p.130)

Wavelet Transforms  MA 043

Chair: Boris Rubin

15:00 – 15:15 Charles N. Moore, Department of Mathematics, Galway, Ireland Distribution function inequalities for sums of wavelets (p.142)

15:15 – 15:30 John Schmeelk, Virginia Commonwealth University, Richmond, Virginia, USA Wavelet Transforms in Generalized Fock Spaces (p.154)

15:30 – 15:45 Boris Rubin, Hebrew University, Jerusalem, Israel Fractional Calculus and Wavelet Transforms in Integral Geometry (p.152)

Short Communications  Thursday, 20 August
Sobolev- and Function Spaces  H 110

Chair: Michal Wojciechowky

16:15 – 16:30 Ritva Hurri-Syrjanen, University of Helsinki, Helsinki, Finland Generalizations of Poincaré inequality (p.131)

16:30 – 16:45 Evgueni M. Semenov, Voronezh State University, Voronezh, Russia Disjoint Systems and Complemented Subspaces in Rearrangement Invariant Spaces (p.155)

16:45 – 17:00 Michal Wojciechowky, Instytut Matematyczny Pan, Warszawa, Poland The multidimensional Marcinkiewicz multiplier theorem for H^{1} spaces on product domains (p.164)

Integral Transforms I  MA 041

Chair: Vakhtang M. Kokilashvili

16:15 – 16:30 Jagdish N. Pandey, Carleton University, Ottawa, Canada The Hilbert Transform of Almost Periodic Functions and Distributions (p.146)

16:30 – 16:45 Alexander N. Meskhi, A. Razmadze Mathematical Institute, Tbilisi, Georgia Solution of some weight problems for the Riemann-Liouville and Weyl operators defined on SHT

(p.141)

16:45 – 17:00 Vakhtang M. Kokilashvili, A. Razmadze Mathematical Institute, Tbilisi, Georgia Solutions of Two-Weight Problems on Boundedness and Compactness for Integral Transforms

(p.135)

Variational Problems  MA 309

Chair: Julian P. Revalski

16:15 – 16:30 Hong-Kun, H.K. Xu, University of Durban-Westville, Durban, South Africa Asymptotic behavior of a gradient flow (p.164)

16:30 – 16:45 Mikhail A. Sytchev, Sobolev Institute of Mathematics, Novosibirsk, Russia Characterization of homogeneous scalar variational problems solvable for all boundary data (p.158)

16:45 – 17:00 Julian P. Revalski, Institute of Mathematics and Informatics, Sofia, Bulgaria Porosity and variational principles (p.152)

Topological Vector Spaces and Algebras  MA 043

Chair: Wieslaw Zelazko

16:15 – 16:30 Aboubakr Bayoumi, King Saud University College of Science, Riyadh, Saudi Arabia Quasi-Differentiability and Applications in Locally Bounded Spaces (p.116)

16:30 – 16:45 Mati Abel, University of Tartu, Tartu, Estonia Structure of locally pseudoconvex algebras (p.112)

16:45 – 17:00 Wieslaw Zelazko, Mathematical Institute, Warszawa, Poland A supersimple algebra (p.166)

Analytic Functions I  H 110

Chair: Žarko Pavićević

17:15 – 17:30 Shigeyoshi Owa, Kinki University, Higashi-Osaka, Japan Some Properties for Convulutions (p.146)

17:30 – 17:45 Yaroslav Ya. Prytula, Lviv State University, Lviv, Ukraine Generalization of Lindelöf Theorem (p.149)

17:45 – 18:00 Žarko Pavićević, Univerzitet Crne Gore, Podgorica, Yugoslavia Meromorphic Functions Generating Normal Families in an Arbitrary Open Subset of the Unit Disk, with respect to any Subgroup of the Group of Conformal Automorphisms of the Unit Disk

(p.147)

Operator Algebras  MA 041

Chair: Huzihiro Araki

17:15 – 17:30 Grigori Amosov, Moscow Institute of Physics and Technology, Dolgopzudni, Russia Cocycle conjugacy classes of Powers’ semiflow of shifts (p.113)

17:30 – 17:45 Alexandr V. Kosyak, Institute of Mathematics, Kiev, Ukraine On von Neumann algebras, generated by regular representations of infinite-dimensional groups

(p.137)

17:45 – 18:00 Huzihiro Araki, Dept. Math., Science University of Tokyo, Kyoto, Japan Factor types of infrared representations of the XY-model (p.114)

Convexity, Semigroups, and Blocking  MA 309

Chair: Karl-Goswin Grosse-Erdmann

17:15 – 17:30 Farhad V. Hüsseinov, Economics Department, Bilkent University, Ankara, Turkey Closedness of convex hull and its applications (p.131)

17:30 – 17:45 Nobuhisa Sakakibara, Faculty of Engineering, Ibaraki University, Hitachi, Japan Perfectness of conelike ∗-semigroups in Q^{k} (p.153)

17:45 – 18:00 Karl-Goswin Grosse-Erdmann, FernUniversität Hagen, Hagen, Germany The blocking technique and the weighted Hardy inequality (p.127)

Orthogonal Polynomials and Wavelets  MA 043

Chair: Thomas C. Kriecherbauer

17:15 – 17:30 Mohsen Taghavi, Shiraz University, College of Sc., Shiraz, Iran Exact values of the autocorrelation coefficients of the Rudin-Shapiro polynomials (p.159)

17:30 – 17:45 Ahmed I Zayed, University of Central Florida, Orlando, Florida, USA Wavelets in Closed Forms (p.166)

17:45 – 18:00 Thomas C. Kriecherbauer, Universität München, München, Germany Asymptotics of Orthogonal Polynomials via Integrable Methods (p.137)

Short Communications  Friday, 21 August
Univalent Functions I  MA 041

Chair: Ruben A. Hidalgo

14:00 – 14:15 Bruce O’Neill, The American University in Cairo, Cairo, Egypt Boundary Behavior of Conformal Mappings Associated with Singular Measures (p.145)

14:15 – 14:30 Teodor Bulboacă, Babeş-Bolyai University, Cluj-Napoca, Romania Differential subordinations and univalent integral operators (p.119)

14:30 – 14:45 Om Parkash Ahuja, Nanyang Technological University, Singapore, Republic of Singapore On Meromorphic Starlike and Convex Functions (p.113)

Integral Transforms II  MA 309

Chair: Kun Soo Chang

14:00 – 14:15 Yukio Kuribayashi, Faculty of Education, Tottori University, Tottori, Japan On pseudofourier transforms (p.138)

14:15 – 14:30 Rama Rawat, Bar-Ilan University, Ramat Gan, Israel Pompeiu problem and L^{p} analogues of Wiener-Tauberian for certain homogeneous spaces (p.151)

14:30 – 14:45 Kun Soo Chang, Yonsei University, Seoul, Korea, Republic of Integral transform and convolution of analytic funtionals on abstract Wiener space (p.120)

Banach Spaces  MA 043

Chair: David Yost

14:00 – 14:15 Parviz Azimi, University of Sistan and Baluchestan, Zahedan, Iran A class of hereditarily complementably ℓ_{p} Banach spaces (p.115)

14:15 – 14:30 Robb C. Fry, University of Northern British Columbia, Prince George, Canada Smooth Bump Functions on Banach Spaces (p.124)

14:30 – 14:45 David Yost, Chulalongkorn University, Bangkok, Thailand Twisted sums of classical Banach spaces (p.165)

Quasiconformal Maps  MA 041

Chair: Bruce O’Neill

15:00 – 15:15 Alexander Jur’evich Vasil’ev, Universidad de Los Andes, Santafe de Bogota, D.C., Colombia Two point distortion theorems for quasiconformal mappings (p.162)

15:15 – 15:30 Evsei M. Dynkin, Dept. of Math., Technion, Haifa, Israel Smoothness of quasiconformal mappings at a point (p.122)

15:30 – 15:45 Vo Dang Thao, University of Ho Chi Minh City, Ho Chi Minh City, Vietnam Area Problem for quasiconformal mappings (p.160)

Sobolev Classes and Matrix Functions  MA 309

Chair: Ilya M. Spitkovsky

15:00 – 15:15 Serguei Vodop’yanov, Sobolev Institute of Mathematics, Novosibirsk, Russia On open and discrete mappings of Sobolev classes with integrable dilatation (p.162)

15:15 – 15:30 Vladimir V. Peller, Department of Mathematics, Manhattan, USA Regularity conditions for vectorial stationary processes (p.147)

15:30 – 15:45 Ilya M. Spitkovsky, The College of William and Mary, Williamsburg, USA Almost periodic factorization and its applications (p.157)

C*-Algebras  MA 043

Chair: George A. Elliott

15:00 – 15:15 Martin Mathieu, Department of Mathematics, Maynooth, Co. Kildare, Ireland A simple local multiplier algebra (p.140)

15:15 – 15:30 Michael Frank, Dept. Mathematics, University of Houston, Houston, Texas, USA Conditional expectations of finite index and discrete group actions (p.124)

15:30 – 15:45 George A. Elliott, University of Toronto, Toronto, Canada An abstract Weyl–von Neumann theorem (p.123)

Short Communications  Saturday, 22 August
Univalent Functions II  H 110

Chair: Arne Stray

16:15 – 16:30 Peter Duren, Department of Mathematics, Ann Arbor, Michigan, USA Hyperbolic Capacity and its Distortion under Conformal Mapping (p.122)

16:30 – 16:45 Ilgiz G. Kayumov, Inst. of Math. and Mech., Kazan University, Kazan, Russia Asymptotical behaviour of logarithmic coefficients of univalent functions (p.134)

16:45 – 17:00 Dmitri V. Prokhorov, Saratov State University, Saratov, Russian Federation Estimates of linear functionals for bounded univalent functions (p.149)

Inequalities and Interpolation  MA 041

Chair: Vladimir I. Ovchinnikov

16:15 – 16:30 Vladislav F. Babenko, Dnepropetrovsk State University, Dnepropetrovsk, Ukraine Inequalities of Landau–Kolmogorov–Nagy Type (p.115)

16:30 – 16:45 Vladimir Alexandrovich Kofanov, Dnepropetrovsk State University, Dnepropetrovsk, Ukraine Exact Inequalities of Kolmogorov Type for Periodic Functions of Small Smoothness (p.135)

16:45 – 17:00 Vladimir I. Ovchinnikov, Voronezh State University, Voronezh, Russia Nonclassical Interpolation Theorems in Spaces of Smooth Functions (p.146)

Measure and Integration  MA 309

Chair: Carmen D. Vlad

16:15 – 16:30 Alexandre V. Mikhalev, Moscow State University, Moscow, Russia Solution of Riesz-Radon problem (p.141)

16:30 – 16:45 Valeriy Konstantinovich Zakharov, Saint Petersburg State University of Technology and Design, Saint Petersburg, Russia A solution of the general Riesz-Radon problem for a Hausdorff space (p.166)

16:45 – 17:00 Carmen D. Vlad, Pace University, New York, NY, USA Outer measures and countably additivity of measures (p.162)

Banach Algebras and Fredholm Maps  MA 043

Chair: Kari E. Ylinen

16:15 – 16:30 Victor G. Zvyagin, Voronezh State University, Voronezh, Russia Topological methods in the theory of perturbations of Fredholm maps and applications (p.167)

16:30 – 16:45 Hans-Olav I. Tylli, Department of Mathematics, University of Helsinki, Finland Rigidity of commutators and elementary operators on Calkin algebras (p.161)

16:45 – 17:00 Kari E. Ylinen, University of Turku, Turku, Finland Positive definite operator functions on groups and related multilinear operators on products of Banach-^{∗}-algebras (p.165)

Hilbert and Krein Spaces  H 110

Chair: Vladimir Strauss

17:15 – 17:30 Elias Flytzanis, Athens University of Economics and Business, Athens, Greece An invariance property of quadratic forms (p.123)

17:30 – 17:45 Michael Gebel, Inst. Angew. Mathem. I, TU Bergakademie, Freiberg, Germany Spectral Theory of Adjointable Operators in Spaces with two Metrics (p.125)

17:45 – 18:00 Vladimir Strauss, Departamento de Matemáticas, U.S.B., Caracas, Venezuela On a model representation for a normal operator in Krein space (p.158)

Multivalued Mappings, Weighted Inequalities, and Means  MA 041

Chair: Lars-Erik Persson

17:15 – 17:30 B.K. Sharma, Pt. Ravishankar Shukla University, Raipur, India A Strong Convergence Theorem for Multivalued Nonexpansive Mappings (p.155)

17:30 – 17:45 Zheng Liu, Anshan Institute of Iron and Steel, Anshan, China, Peoples Republic of Inequalities for Some Means (p.139)

17:45 – 18:00 Lars-Erik Persson, Luleå University, Luleå, Sweden Sharp multidimensional multiplicative inequalities for weighted Lebegue spaces (p.147)

Dunkl Operators and Hyperfunctions  MA 309

Chair: Dohan Kim

17:15 – 17:30 Margit Rösler, Technische Universität München, München, Germany Positivity of Dunkl’s intertwining operator (p.152)

17:30 – 17:45 Stevan Pilipović, University of Novi Sad, Novi Sad, Yugoslavia Suppleness of Some Quotient Spaces of Microfunctions (p.147)

17:45 – 18:00 Dohan Kim, Dept. of Math., Seoul National Univ., Seoul, Korea, Republic of Moment problem for hyperfunctions and the heat kernel method (p.134)

Measure Problems  MA 043

Chair: Jiří Jarník

17:15 – 17:30 Raymond J. Grinnell, University of the West Indies, Bridgetown, Barbados Two real functions associated with convolutions of measures (p.127)

17:30 – 17:45 Daniar H. Mushtari, Inst. of Math. and Mech., Kazan University, Kazan, Russia Orthomodular posets of projections on linear spaces and signed measures on them (p.143)

17:45 – 18:00 Jiří Jarník, Pedagogicka Fakulta UK, Praha, Czech Republic Convergence theorem for Henstock–Kurzweil integral based on the convergence of primitives (p.132)

Short Communications  Monday, 24 August
Several Complex Variables  MA 041

Chair: Stefan B.A. Jakobsson

14:00 – 14:15 Yoshihiro Abe, Toyama University, Yokohama, Japan Universal functions in several variables (p.112)

14:15 – 14:30 Elias Wegert, TU Bergakademie Freiberg, Freiberg, Germany Analytic discs attached to totally real surfaces (p.163)

14:30 – 14:45 Gabriela Kohr, Babeş-Bolyai University, Cluj-Napoca, Romania Univalent mappings of several complex variables (p.135)

Real Analysis  MA 309

Chair: Gilbert Helmberg

14:00 – 14:15 Jan Hejcman, University of Education, Hradec Králové, Praha, Czech Republic A description of quasiconvex functions (p.129)

14:15 – 14:30 Attila Gilányi, Lajos Kossuth University, Debrecen, Hungary On polynomial functions, monomial functions, and the Dinghas interval derivative (p.125)

14:30 – 14:45 Joris van der Hoeven, École Polytechnique, Paris, France Proof of a conjecture of Hardy about L-functions (p.161)

Approximation Theory  MA 043

Chair: Miguel A. Jimenez

14:00 – 14:15 Alexander Yakovlevich Radyno, Belarusian State University, Minsk, Belarus The degree of rational approximation of a holomorphic function (p.149)

14:15 – 14:30 Yuri Kryakin, Odessa State Polytechnic University, Odessa, Ukraine Polynomial Interpolation and Whitney’s Constants (p.138)

14:30 – 14:45 Miguel A. Jimenez, BUAP Puebla, Puebla, Mexico The Degree of Best Approximation in the Lipschitz Norm by Trigonometric Polynomials (p.133)

Harmonic Functions  MA 041

Chair: Elias Wegert

15:00 – 15:15 Anna Łazińska, University of Łódź, Łódź, Poland On a class of mappings harmonic in the half-plane, with a normalization at infinity (p.138)

15:15 – 15:30 Promarz M. Tamrazov, Inst. of Math. of National Acad. of Sci. of Ukraine, Kiev, Ukraine The solution of Eremenko’s extremal problem on harmonic functions (p.159)

15:30 – 15:45 Eleutherius E. Symeonidis, Katholische Universität Eichstätt, Eichstätt, Germany The Poisson integral as an appropriate mean value (p.159)

A New View of Classical Phenomena  MA 309

Chair: Gilbert Helmberg

15:00 – 15:15 Andreas M. Hinz, Technische Universität München, München, Germany Catenaria vera — the true catenary (p.130)

15:15 – 15:30 J. Marshall Ash, DePaul University, Chicago, USA Nonuniqueness for smooth uniformly bounded orthonormal systems (p.114)

15:30 – 15:45 Gilbert Helmberg, Institut f. Math. u. Geometrie, Universität Innsbruck, Innsbruck, Austria A corner-point Gibbs phenomenon for Fourier series in two dimensions (p.129)

Vector Lattices  MA 043

Chair: Jacobus J. Grobler

15:00 – 15:15 Rodica Mihaela Danet, Technical University of Bucharest, Bucharest, Romania The extension of some linear operators (p.121)

15:15 – 15:30 Nicolae Danet, Technical University of Bucharest, Bucharest, Romania Subspaces in directed vector spaces (p.121)

15:30 – 15:45 Jacobus J. Grobler, Potchefstroom University for CHE, Potchefstroom, South Africa The polar decomposition of disjointness preserving operators (p.127)

Short Communications  Tuesday, 25 August
Miscellaneous  MA 309

Chair: Eberhard Kirchberg

15:00 – 15:15 Eberhard Kirchberg, Humboldt–Universität, Berlin, Germany Classification of non-simple nuclear C*-algebras

15:15 – 15:30 Volodymyr K. Maslyuchenko, Chernivtsi University, Chernivtsi, Ukraine Separately continuous mappings: Van Vleck spaces and inverse problems

Hilbert Spaces, Wavelets, and Fractals  H 110

Chair: Gerald S. Goodman

16:15 – 16:30 Kofi V.S. Assiamoua, Université du Bénin, Lomé, Togo A product of sesquilinear mappings on Hilbert spaces (p.114)

16:30 – 16:45 Raquel-Garcia Catalan, University of Zaragoza, Zaragoza, Spain From MRA to Tight Frames (p.120)

16:45 – 17:00 Gerald S. Goodman, University of Florence, Florence, Italy Dilation equations and Hutchinson fractals (p.126)

Real and Complex Approximation  MA 041

Chair: Mirela Kohr

16:15 – 16:30 Nikolai Ya. Radyno, Institute of Mathematics, Minsk, Belarus Some elements of a General Spectral Theory of Banach modules and their applications (p.150)

16:30 – 16:45 Oxana V. Motornaya, Dnepropetrovsk State University, Dnepropetrovsk, Ukraine On the Approximation of the Functions by Algebraic Polynomials (p.142)

16:45 – 17:00 Andrey Olegovich Matveev, Institute of Mathematics and Mechanics, Ekaterinburg, Russia Error estimates for scattered data interpolation

Spectral Analysis and Operator Algebras  MA 309

Chair: Sa Ge Lee

16:15 – 16:30 Antoine-A.P. Bredimas, University Paris 7, Paris, France Schlicht functions and exotic sprectral analysis: Applications to some conjectures and problems in analysis (p.118)

16:30 – 16:45 Valentyn Ja. Golodets, Institute for Low Temperature Physics, Kharkov, Ukraine Dynamical entropy and quantum K-systems (p.126)

16:45 – 17:00 Sa Ge Lee, Seoul National University, Seoul, Korea, Republic of The generalized normal state space and unital normal completely positive map (p.139)

Matrix and Operator Functions  MA 043

Chair: George N. Khimshiashvili

16:15 – 16:30 Marcin J. Zygmunt, Institute of Mathematics PAN, Krakow, Poland Matrix Moment Problem (p.167)

16:30 – 16:45 Frank-Olme Speck, Universidade Técnica de Lisboa, Lisboa, Portugal Relations between convolution type operators on intervals and on the half-line (p.157)

16:45 – 17:00 George N. Khimshiashvili, University of Łódź, Łódź, Poland Homotopy classes of elliptic boundary problems (p.134)

Analytic Functions II  H 110

Chair: Mohamed Abdel-Megied

17:15 – 17:30 Gueorgui S. Litvinchuk, Universidade da Madeira, Funchal, Portugal On solvability theory of Hilbert boundary value problem with a Carleman shift (p.139)

17:30 – 17:45 Mirela Kohr, Babeş-Bolyai University, Cluj-Napoca, Romania A direct boundary integral formulation for a Stokes flow problem (p.135)

17:45 – 18:00 Rauno Aulaskari, University of Joensuu, Joensuu, Finland Q_{p}-spaces on Riemann surfaces (p.115)

Nonlinear Operators  MA 041

Chair: Richard M. Dudley

17:15 – 17:30 Radoje Šćepanović, Univerzitet Crne Gore, Podgorica, Yugoslavia On Solvability of Some Nonlinear Equations in Hilbert Spaces (p.154)

17:30 – 17:45 Vladilen A. Trenogin, Moscow State Steel & Alloys Institute, Moscow, Russia Parameter continuation method and solvability of nonlinear equations (p.160)

17:45 – 18:00 Richard M. Dudley, MIT, Cambridge, MA, USA Differentiability of six nonlinear operators on nonsmooth functions (p.122)

Convexity, Banach Spaces, and Isometries  MA 309

Chair: Themistocles M. Rassias

17:15 – 17:30 Alexander Ya. Kruger, International Institute of Labour and Social Relations, Minsk, Belarus On extremality of set systems (p.137)

17:30 – 17:45 Ryszard Grza̧ślewicz, Wrocław University of Technology, Wrocław, Poland Quasi-interior points for convex sets (p.128)

17:45 – 18:00 Themistocles M. Rassias, National Technical University of Athens, Athens, Greece On the Mazur-Ulam Theorem and the Aleksandrov Problem of Conservative Distances (p.151)

Sequence and Function Spaces  MA 043

Chair: Jan H. Fourie

17:15 – 17:30 Abdul J. Badiozzaman, Chamran University, Ahvaz, Iran On a Tauberian Theorem of G.M. Petersen for Abel Summability (p.116)

17:30 – 17:45 Detlef Seydel, FH-Braunschweig/Wolfenbüttel, Wolfsburg, Germany Matrix Maps of Quasiconvex Null Sequences (p.155)

17:45 – 18:00 Jan H. Fourie, Potchefstroom University, Potchefstroom, South Africa On convergence of sections of weakly p-summable sequences (p.123)

Short Communications  Wednesday, 26 August
Special Functions  MA 041

Chair: Promarz M. Tamrazov

14:00 – 14:15 Raoul R.F.G. Gloden, Cours Universitaires, Luxembourg, Luxembourg Propriétés des polynômes orthogonaux. Développements de cas particuliers (p.126)

14:15 – 14:30 Kathy A. Driver, University of Witwatersrand, Johannesburg, South Africa Zeros of Hypergeometric Functions (p.121)

14:30 – 14:45 Momčilo Bjelica, University of Novi Sad, Zrenjanin, Yugoslavia Asymptotic of means (p.117)

Metrizable Spaces, Riesz Spaces, and Integral Representations  MA 309

Chair: Paul Ressel

14:00 – 14:15 Shozo Koshi, Hokkaido Institute of Technology, Sapporo, Japan Partially ordered linear space (p.136)

14:15 – 14:30 Luis M. Sánchez Ruiz, Universidad Politécnica de Valencia, Valencia, Spain On Metrizable Barreled Spaces, Cardinals and Dimensionality (p.153)

14:30 – 14:45 Paul Ressel, Kath. Universität Eichstätt, Eichstätt, Germany Semigroup representations, positive definite functions, and abelian C*-algebras (p.152)

Operator Algebras and Systems  MA 043

Chair: Palle E.T. Jorgensen

14:00 – 14:15 B.V. Rajarama Bhat, Indian Statistical Institute, Bangalore, India Minimal dilation of quantum dynamical semigroups to semigroups of endomorphisms (p.116)

14:15 – 14:30 Wend Werner, Universität Paderborn, Paderborn, Germany Operator Systems without Unit (p.163)

14:30 – 14:45 Palle E.T. Jorgensen, The University of Iowa, Iowa City, IA, USA Ruelle operators: Functions which are harmonic with respect to a transfer operator (p.133)

Kleinian Groups  MA 041

Chair: Alexander Jur’evich Vasil’ev

15:00 – 15:15 Ruben A. Hidalgo, UTFSM, Valparaiso Casilla, Chile A_{4}, A_{5}, and S_{4} of Schottky Type (p.130)

15:15 – 15:30 Weiqi Yang, Beijing Institute of Technology, Beijing, P.R. of China Discrete groups and dense groups (p.164)

15:30 – 15:45 Javier Pérez, Universidad Nacional de Educación a Distancia, Madrid, Spain Harmonic differentials on Klein Surfaces (p.147)

Analytic Functions and Toeplitz Operators  MA 309

Chair: Marek Ptak

15:00 – 15:15 Ali Abkar, Lund University, Lund, Sweden A comparison theorem concerning hyperinvariant subspaces on finitely connected domains (p.113)

15:15 – 15:30 Slavko Simic, Belgrade, Yugoslavia Entire Functions of Finite Order and Regularly Varying Sequences (p.156)

15:30 – 15:45 Marek Ptak, University of Agriculture, Kraków, Poland Toeplitz operators; dichotomy of reflexivity and transitivity (p.149)

Operator Algebras and Algebras of Operators  MA 043

Chair: Vladimir Manuilov

15:00 – 15:15 Raphael S. Ponge, Univ. Paris-Sud, Dpt. of Math., Orsay, France Wodzicki Residue for Heisenberg Manifolds (p.148)

15:15 – 15:30 Georgios F. Nassopoulos, University of Athens, Athens, Greece Coproducts and projective presentations of operator algebras (p.144)

15:30 – 15:45 Vladimir Manuilov, Moscow State University, Moscow, Russia Almost representations of discrete groups and K-theory (p.140)

Poster Session  Wednesday, 19 August 14:00 – 15:45  MA 144

Aldric L. Brown, The Pnajab University, Chandigarh, India

Operator convex functions of two variables (p.118) Mao-Ting Chien, Soochow University, Taipei, Taiwan

A parametrization of the c-numerical ranges of certain matrices (p.120)

William C. Connett, University of Missouri, St. Louis, MO, USA Measure algebras that have oblate spheroidal wave functions as characters (p.121)

Sonja Mouton, University of Stellenbosch, Matieland, South Africa Some spectral properties in ordered Banach algebras (p.143)

Marjatta Näätänen, University of Helsinki, Helsinki, Finland Arithmetic Fuchsian groups of signature (0; a, b, c, oo) (p.143)

Samyon R. Nasyrov, Kazan State University, Kazan, Russia Robin capacity and lift of infinitely thin airfoils (p.144)

David L. Salinger, University of Leeds, Leeds, Great Britain Lifting the Difference Spectrum (p.153)

Ajit Iqbal Singh, Ohio University, Athens, Ohio, USA Maximal quotient rings of algebras of functions and operators (p.156)

Arne Stray, University of Bergen, Kristiansand, Norway Polynomial approximation of analytic functions with finite Dirichlet integral (p.158)

Chia-Chi Tung, Mankato State University, Mankato, MN, USA A Note on the Hilbert Nullstellen Satz (p.161)

Dmitri Zaitsev, Universität Tübingen, Tübingen, Germany On holomorphic boundary extensions, web theory, and invariants of Grassmanians (p.165)

Jaroslav Zemánek, Polish Academy of Sciences, Warszawa, Poland Powers and resolvents (p.167)

Poster Session  Friday, 21 August 14:00 – 15:45  MA 144

Jelena Ausekle, Tartu University, Tartu, Estonia

Compactness of Operators Acting in the Lorentz and Orlicz Sequence Spaces (p.115) Prodipeswar Bhattacharyya, Indian Institute of Technology, Chennai, India

On the domain of normality in iteration theory (p.117) Davor Butkovic, University of Zagreb, Zagreb, Croatia

Transformations of regular almost convergent sequences (p.119) Rainis Haller, Tartu University, Tartu, Estonia

M(r, s)-inequality and reflexivity (p.128) Katsunori Iwasaki, Kyushu University, Fukuoka, Japan

Polyhedral harmonic functions and invariant theory (p.132) Volodymyr M. Kadets, Kharkov State University, Kharkov, Ukraine

The Lyapunov property of certain Banach spaces (p.133) Shuichi Ohno, Nippon Institute of Technology, Miyashiro Saitama, Japan

Topological structure of the space of composition operators on H^{∞} (p.145) Eve Oja, Tartu University, Tartu, Estonia

Lifting approximation properties and ideals of compact and approximable operators (p.145) Zdzislaw Otachel, Inst. of Appl. Math. Agric. Univ. of Lublin, Lublin, Poland

Cone Preorders in Measurable Function Spaces (p.145) Märt Põldvere, Tartu University, Tartu, Estonia

On subspaces of Banach spaces having Phelps’ uniqueness property (p.148) Sen-Yen Shaw, National Central University, Chung-Li, China, Taiwan

Convergence rates of regularized approximation processes (p.156) Panagiotis D. Stratigos, Long Island University, Brooklyn, N.Y., USA

On general measure repleteness and completeness (p.158)

Poster Session  Wednesday, 26 August 14:00 – 15:45  MA 144

Ants Aasma, Tallinn Pedagogical University, Tallinn, Estonia Summability with speed and comparison of orders of approximation of Fourier expansions by different matrix methods of summability (p.112)

El Hassan E.H. El Kinani, Faculté des Sciences Mohammed V, Errachidia, Morocco

q-Derivation, Sine algebra, and Ut(sl(2)) (p.123) Stefan B.A. Jakobsson, University of Lund, Lund, Sweden

Applications of weighted Hele-Shaw flow and kernel function techniques to Green functions for weighted biharmonic operators (p.132)

Volodymyr M. Kolodyazhniy, Kharkiv Aviation Institute, Kharkiv, Ukraine

Nonlinear Identification Problem Solution Using the Wavelets of the Atomic Functions (p.136) Hrvoje Kraljevic, University of Zagreb, Zagreb, Croatia

Landau’s type inequalities for contraction cosine functions (p.137) Carlos Lizama, Universidad de Santiago de Chile, Santiago Correo, Chile

A Characterization of Uniform Continuity for Volterra Equations in Hilbert Spaces (p.140) Igor-Ya. Novikov, Voronezh State University, Voronezh, Russia

Time-frequency localization of modified Daubechies wavelets (p.144) Alan L. Schwartz, University of Missouri-St. Louis, St. Louis, MO, USA

Hypergroups and their maximum subgroup (p.154) Katsunori Shimomura, Ibaraki University, Mito, Japan

On the transformations which preserve polytemperatures (p.156) Feng-Yu Wang, Beijing Normal University, Beijing, P.R. China

Harnack inequality for heat semigroups and log-Sobolev inequalities (p.163) Oxana A. Ziza, Moscow State Institute of Electronics and Mathematics, Moscow, Russia

On absolute summability of orthogonal series (p.167)